### v1

# Iterative Partition Search Variational Quantum Algorithm for Solving Shortest Vector Problem

Abstract
The Partition Search Algorithm (PSA) and the Iterative Quantum Optimization with an Adaptive Problem (IQOAP) framework are two existing Variational Quantum Algorithms (VQAs) for solving the Shortest Vector Problem (SVP), but both suffer from certain limitations. In this work, we proposed the Iterative Partition Search Algorithm (IPSA), which is a targeted synthesis and refinement of these preceding methods. Our algorithm inherits the core idea of “partitioning to circumvent the zero vector” from PSA and the “iterative lattice basis reduction” framework from IQOAP. A key feature of IPSA is the “1-tailed search spaces”, which can be viewed as a highly constrained variant of PSA’s partitioning strategy, specifically designed for optimal performance within IQOAP’s iterative structure. We supplant IQOAP’s fixed iteration count with a dynamic, stack-managed process and substitute a more expressive and shallower circuit structure for its original ansatz. Crucially, the 1-tailed design fundamentally ensures that every successful VQA execution yields an effective lattice basis update, thereby eliminating the issue of ineffective iterations in IQOAP. This evolutionary path of refinement allows IPSA to overcome the drawbacks of its predecessors precisely. Numerical simulations on 4- to 6-dimensional SVP instances demonstrate that IPSA achieves at least a 73% improvement in success rate for finding optimal solutions and over a 35% improvement in average solution quality compared with the methods above while maintaining comparable total circuit depth.
I  Introduction
The Shortest Vector Problem (SVP) is an NP-hard problem in lattice theory, underpinning the security of numerous post-quantum cryptographic schemes. Lattice is defined as a discrete additive subgroup of ${\mathbb{R}}^{n}$, typically generated by a basis matrix. These basis vectors form a lattice through their integer linear combinations. SVP seeks to identify the shortest non-zero vector within a lattice structure. Classical algorithms such as enumeration [1, 2, 3] and sieving [4, 5, 6] can solve SVP but require substantial computational costs. Quantum algorithms offer promising alternatives by exploiting Grover’s search [7] and quantum walks [8, 9, 10, 11] to achieve significant speedups. Recent quantum advances have reduced sieving complexity to $2^{0.2563n+o(n)}$ [10] and demonstrated quadratic speedups for enumeration method [11].
However, these quantum advantages typically require fault-tolerant quantum computers, which remain beyond current technological capabilities. Present-day Noisy Intermediate-Scale Quantum (NISQ) devices are characterized by limited qubit counts and inherent noise, motivating the development of Variational Quantum Algorithms (VQAs) [12, 13, 14, 15, 16]. VQA represents hybrid quantum-classical computing paradigms that employ Parameterized Quantum Circuits (PQC), described by parameters $\bm{\theta}$, to prepare trial states $|\psi(\bm{\theta})\rangle$. A classical optimizer iteratively adjusts $\bm{\theta}$ to minimize a cost function $C(\bm{\theta})=\langle\psi(\bm{\theta})|H|\psi(\bm{\theta})\rangle$, where $H$ represents a Hamiltonian encoding the optimization problem. With a shallower circuit and inherent noise tolerance [17], VQA is well-suited for NISQ devices, with prominent examples ranging from foundational algorithms like the Variational Quantum Eigensolver (VQE) [18] and the Quantum Approximate Optimization Algorithm (QAOA) [19], to broader applications in areas like quantum federated learning [20] and quantum neural networks [21, 22, 23].
Investigating VQA-based solvers for SVP is a meaningful endeavor, as it helps assess the potential threat that NISQ devices pose to post-quantum cryptography. Applying VQA to SVP involves three key aspects of consideration, and several existing studies have begun to address each of these points.
First, mapping the infinite coefficients search space $\mathbb{Z}^{n}$ of SVP onto finite qubit count requires binary encoding strategy [24] and the establishment of bounded search ranges [25, 26], where the range boundaries are typically guided by classical preprocessing algorithms such as Lenstra–Lenstra–Lovász (LLL) [27] and Hermite-Korkin-Zolotarev (HKZ) [1]. Recently, Zhu, Joseph et al. proposed the Iterative Quantum Optimization with an Adaptive Problem (IQOAP) framework for solving SVP [28], achieving qubit requirement reduction compared with conventional boundary constraint methods [25, 26].
Second, the straightforward Hamiltonian construction for the vector norm objective leads to a trivial all-zero ground state, necessitating sophisticated approaches to avoid this trivial solution. Standard avoidance strategies, such as introducing penalty terms and redesigning the cost function, come with certain costs due to the difficulty in tuning penalty coefficients and the requirement for additional qubits [25, 29, 30]. The Partition Search Algorithm (PSA) [31], originally proposed and validated in adiabatic quantum computing, offers an elegant approach to bypass this trivial solution issue without incurring any additional costs. Although initially developed within the adiabatic quantum computing framework, PSA demonstrates equivalent efficacy in VQA for avoiding trivial solutions.
Third, the design of PQC plays a crucial role in algorithm effectiveness. Standard ansatz like QAOA, while effective for certain optimization problems [32, 33], typically require deep circuit architectures for SVP due to the all-to-all connectivity inherent in the problem Hamiltonian [28, 30, 26]. This limitation motivates the adoption of alternative circuit structures, such as the Hardware-Efficient Ansatz (HEA) [34, 35, 36, 37, 38], which can provide strong expressive power with shallower circuits. Recently, we have also noted other mitigation strategies, such as using Fixed-Angle QAOA variants [39, 40, 41] or leveraging classical post-processing [42].
However, both PSA and IQOAP, despite their contributions, exhibit significant room for improvement. The complex optimization landscape of PSA often leads to convergence on low-quality suboptimal solutions, while its qubit requirement can be further reduced. On the other hand, IQOAP does not inherently avoid the trivial all-zero solution, and its effectiveness is constrained by a weak circuit, a fixed number of iterations, and the potential for ineffective iterations that waste computational resources.
To address these issues, we propose the Iterative Partition Search Algorithm (IPSA), which is designed as a targeted synthesis and refinement of these two preceding methods. IPSA inherits the core idea of “partitioning to circumvent the zero vector” from PSA and the “iterative lattice basis reduction” framework from IQOAP. A key feature of IPSA is the “1-tailed search space”, which acts as a highly constrained variant of PSA’s strategy, specifically engineered for optimal performance within this iterative structure. We replace the fixed iteration count with a dynamic, stack-managed process and substitute the initial ansatz with a more expressive circuit structure. Crucially, this 1-tailed design ensures that every successful VQA execution yields an effective lattice basis update, thereby fundamentally solving the issue of ineffective iterations. These refinements enable IPSA to achieve superior success rates and solution quality without a significant increase in computational cost overhead, thus advancing the application of VQA for solving SVP.
The remainder of this paper is organized as follows. In Section II, we review two existing methods, the PSA and the IQOAP. In Section III, we present our proposed IPSA and detail its key components, including the 1-tailed search spaces, the stack-based management of search partitions, and the improved quantum circuit design. Section IV reports the results of our numerical simulations, where we benchmark the three algorithms on two distinct sets of SVP instances. In Section V, we provide our conclusions and a brief outlook for future applications.
II  Foundation: Partition Search and Iterative Optimization Framework
The PSA [31] addresses the zero-vector issue without incurring additional computational overhead. It partitions the coefficient space of basis vector combinations into non-overlapping regions and solves the SVP independently within each region. For an $n$-dimensional lattice, the coefficient space is divided into regions $X_{1},\ldots,X_{n}$, where each region $X_{i}$ is defined as:
$X_{i}=\{(x_{1},\dots,x_{i},0,\dots,0)^{T}\in\mathbb{Z}^{n}:x_{i}\geq 1\}$ (1)
This partitioning ensures that regions are mutually exclusive and none contains the zero vector, thereby eliminating the trivial solution.
To construct the problem Hamiltonian for each region $X_{i}$, we assume each coefficient is represented using binary encoding with appropriate qubit allocation. The first i1i-1 coefficients are encoded as:
$\hat{x_{r}}=\frac{1}{2}-\sum_{j=0}^{m}2^{j-1}\sigma^{z}_{r,j}$ (2)
representing integers in the range $(-2^{m},2^{m}]$, where $\sigma^{z}_{r,j}$ denotes the Pauli-Z operator acting on the $j$-th qubit of the $r$-th coefficient. The $i$-th coefficient, constrained by $x_{i}\geq 1$, is encoded as:
$\hat{x_{i}}=\sum_{j=0}^{m-1}2^{j-1}(\sigma^{z}_{i,j}+1)+1$ (3)
representing integers in the range $1,2^{m}]$. The Hamiltonian for region $X_{i}$ is constructed as the squared Euclidean norm:
$H_{i}=\sum_{k=1}^{n}\left(\sum_{r=1}^{i-1}\hat{x_{r}}b_{r,k}+\hat{x_{i}}b_{i,k}\right)^{2}$ (4)
where $b_{r,k}$ denotes the $k$-th component of the $r$-th lattice basis vector.
The ground state of each $H_{i}$ yields the shortest vector within the corresponding partition, and the global minimum among all partitions provides the SVP solution. The total qubit requirement $M$ depends on the sum of bits needed to represent each coefficient $m$, which is determined by the preprocessing of the lattice basis. The analysis in Ref. [25] shows that for an HKZ-reduced basis, the requirement is $M=\mathcal{O}(n\log n)$, while Ref. [26] concludes that using an LLL-reduced basis requires $M=n(n+1)$ qubits. Although PSA effectively circumvents the trivial solution, its complex optimization landscape poses a significant risk of convergence on low-quality suboptimal solutions. Moreover, its implementation via boundary constraint methods still requires a substantial number of qubits, presenting opportunities for further qubit count reduction.
To reduce the qubit utilization in boundary constraint methods, the IQOAP framework [28] performs VQA optimization within a smaller, fixed coefficient range, leveraging the fact that even suboptimal solutions can progressively reduce the lattice basis through iterative refinement. For example, in 4-dimensional SVP instances, each $x_{r}$ is encoded using only 22 qubits via an encoding like:
$\hat{x_{r}^{\prime}}=\frac{1-\sigma^{z}_{r,0}-2\sigma^{z}_{r,1}}{2}$ (5)
Correspondingly, the Hamiltonian expression is:
$~H^{\prime}=\sum_{k=1}^{n}(\sum_{r=1}^{n}\hat{x_{r}^{\prime}}b_{r,k})^{2}$ (6)
The IQOAP framework improves the basis quality through an iterative refinement process, which is outlined in Algorithm 1. In each iteration, if the VQA finds a vector $\bm{v}$ shorter than a current basis vector $\bm{b}_{j}$, then $\bm{b}_{j}$ is replaced by $\bm{v}$ (provided the lattice remains unchanged). Following the original study, this process is repeated for a fixed number of times. For instance, the authors state that 50 iterations suffice to solve 4-dimensional SVP instances with a high probability [28].
Algorithm 1 IQOAP Framework
1:Input: Basis $B=[\bm{b}_{1},\dots,\bm{b}_{n}]$; initial counter $iter=0$.
[…]
3:Solve SVP with QAOA ansatz, obtaining the result $\bm{v}$.
[…]
5:the lattice remains unchanged.
6:If there are multiple vectors that can be replaced in
7:Step 3, choose the longest one.
[…]
While PSA can potentially benefit from IQOAP’s qubit reduction approach, IQOAP itself suffers from several critical limitations. First, it employs a fixed iteration count and single-layer QAOA ansatz with constrained parameters, which severely restricts the quantum circuit’s expressiveness for higher-dimensional SVP instances. Second, if the vector $\bm{v}$ obtained in Step 2 cannot replace any basis vector $\bm{b}_{j}$ without altering the lattice structure, the entire iteration becomes a wasted computation.
To address these limitations while preserving the advantages of both approaches, we propose the IPSA, which combines PSA’s partition strategy with an advanced iterative framework. We employ a stack data structure to iterate across different partitions dynamically, addressing IQOAP’s fixed iteration count limitation. Additionally, we replace IQOAP’s QAOA ansatz with the HEA, which offers greater expressiveness with reduced circuit depth. Furthermore, our proposed 1-tailed search spaces technique not only achieves greater qubit reduction and improved solution quality compared with IQOAP but also eliminates wasted computations that may occur in IQOAP.
III  Iterative Partition Search Algorithm
In this section, we present the overall framework of the IPSA, including the specific design of the parameterized quantum circuit and the detailed analysis of 1-tailed search spaces.
IPSA operates on an iterative principle, using a stack to manage the search partitions dynamically. This stack-based strategy is designed to prioritize re-solving SVP within smaller partitions immediately after a basis vector is updated. The goal of this strategy is to ensure the basis is well-reduced (i.e., composed of shorter, more orthogonal vectors) before proceeding to larger partitions, thereby aiming to reduce the overall computational cost. The complete algorithm is outlined in Algorithm 2.
Algorithm 2 Iterative Partition Search Algorithm
1:Input: Basis $B=[\bm{b}_{1},\dots,\bm{b}_{n}]$; initial empty stack $S$.
2:Output: $\bm{b}_{1}$.
3:Sort $B$ by increasing vector norms.
[…]
5:Pop partition $Y_{i}$ from $S$.
[…]
7:Replace $\bm{b}_{i}$ with $\bm{v}$ and keep $B$ ordered.
8:Push partitions $Y_{i},\cdots,Y_{r}$ onto $S$ in sequence,
[…]
The 1-tailed search spaces $Y_{i}$ defined for an $n$-dimensional lattice as:
$~Y_{i}$ $=$ $\{(y_{1},\cdots,y_{i-1},1,0,\cdots,0)^{T}\in\mathbb{Z}^{n}\}$ (7)
Unlike PSA’s partitions $X_{i}$ in Eq. 1, the 1-tailed spaces $Y_{i}$ fix the $i$-th coefficient to 1. When solving SVP within partition $Y_{i}$, the resulting solution vector is
$~\bm{v}=\sum_{j=1}^{i-1}y_{j}\bm{b}_{j}+\bm{b}_{i}$ (8)
(a)  HEA
(b)  QAOA Ansatz
Figure 1:  Comparison of PQC architectures. (a) HEA: one layer of RY rotations followed by CNOT entanglement and another RY layer. (b) QAOA: alternating objective and mixing layers. For the SVP application, the objective layer is implemented with all-to-all $Z$-interactions, resulting in a deep circuit.
This design, which fixes the coefficient of the $i$-th basis vector to 1, offers several crucial advantages. It reduces qubit requirements by eliminating the encoding for one coefficient. It also simplifies the lattice basis update procedure. Since the $i$-th coefficient is fixed to 1, any solution $\bm{v}$ found in $Y_{i}$ can directly replace $\bm{b}_{i}$ if $\|\bm{v}\|<\|\bm{b}_{i}\|$ while preserving the lattice structure, as formalized in Theorem 1. Most critically, this constrained partitioning simplifies the cost function landscape, mitigating interference from other near-optimal vectors of very similar length. Consequently, the VQA’s search for the true optimum becomes more focused, which significantly increases the probability of convergence and improves the overall solution quality. The efficacy of the 1-tailed search spaces effect is experimentally validated in Section IV.4.
Then, by integrating these focused search steps into an iterative framework, we define the core of our 1-tailed partition strategy. This strategy allows the algorithm to collect multiple distinct candidate solutions across its iterations. The final identification of the shortest vector is then made through a deterministic comparison of these candidates, which enhances the overall probability of success.
Theorem 1
Let $B=[\bm{b}_{1},\dots,\bm{b}_{n}]$ be a basis for a lattice $\mathcal{L}$. If a vector $\bm{v}\in\mathcal{L}$ can be expressed as $\bm{v}=\sum_{j=1}^{n}c_{j}\bm{b}_{j}$ with $c_{j}\in\mathbb{Z}$, and for some $k\in[1,n]$, $|c_{k}|=1$, then the set of vectors $B^{\prime}=[\bm{b}_{1},\dots,\bm{b}_{k-1},\bm{v},\bm{b}_{k+1},\dots,\bm{b}_{n}]$ also forms a basis for $\mathcal{L}$.
Proof. Since $|c_{k}|=1$, we can express $\bm{b}_{k}$ as $\bm{b}_{k}=c_{k}^{-1}(\bm{v}-\sum_{j\neq k}c_{j}\bm{b}_{j})$. As $c_{k}^{-1}=\pm 1$, $\bm{b}_{k}$ is an integer linear combination of the vectors in $B^{\prime}$. Thus, $\mathcal{L}(B)\subseteq\mathcal{L}(B^{\prime})$. Since $\bm{v}\in\mathcal{L}(B)$, all vectors in $B^{\prime}$ are in $\mathcal{L}(B)$, so $\mathcal{L}(B^{\prime})\subseteq\mathcal{L}(B)$. Therefore, $\mathcal{L}(B^{\prime})=\mathcal{L}(B)$, and $B^{\prime}$ is a basis for $\mathcal{L}$. $\square$
Since the primary focus of our work is not on developing a novel PQC, we adopt a HEA architecture similar to that used in Ref. [25]. This choice, over the more common QAOA ansatz, is motivated by the structure of the SVP Hamiltonian. The SVP Hamiltonian (e.g., Eq. 10) features all-to-all interactions among the qubits. For QAOA ansatz, this inherent connectivity translates into a deep and dense objective layer, as shown in Fig. 1(b), increasing its susceptibility to errors on NISQ devices. In contrast, the HEA architecture, illustrated in Fig. 1(a), offers intense expressivity with a typically shallower circuit depth, making it more suitable for navigating the complex energy landscape of SVP. While HEA can be susceptible to barren plateaus, this risk is substantially mitigated in our framework by the reduced 1-tailed search spaces of each VQA instance and our use of a gradient-free classical optimizer. The efficacy of the PQC effect is experimentally validated in Section IV.2.
For coefficient encoding in partition $Y_{i}$, the $i$-th coefficient is fixed at 1, requiring no qubits. The remaining i1i-1 variable coefficients $y_{r}$ (for $r<i$) are each encoded using a specific number of qubits. For the qubit allocation per variable coefficient, we adopt the approach based on experimental observations and related studies [25, 28, 26]. We find that for an $n$-dimensional lattice, representing each variable coefficient $y_{r}$ with $\lfloor\log n\rfloor$ qubits is typically sufficient for IPSA to identify the SVP solution with high probability. Therefore, an appropriate encoding for these variable coefficients is:
$\hat{y_{r}}=\frac{1}{2}-\sum_{j=0}^{\lfloor\log n\rfloor-1}2^{j-1}\sigma^{z}_{r,j}$ (9)
This encoding represents integers in the range $({-2}^{\lfloor\log n\rfloor-1},2^{\lfloor\log n\rfloor-1}]$. The Hamiltonian for partition $Y_{i}$ is then given by:
$H_{i}^{\prime}=\sum_{k=1}^{n}\left(\sum_{r=1}^{i-1}\hat{y_{r}}b_{r,k}+b_{i,k}\right)^{2}$ (10)
Figure 2:  Qubit requirements of IPSA versus the PSA with different classical preprocessing methods. The plot compares IPSA (blue) with PSA using LLL-reduced bases (red) and HKZ-reduced bases (green). While its qubit scaling is comparable to PSA with HKZ, IPSA avoids an exponential-time HKZ preprocessing routine and maintains a significant advantage over PSA with LLL.
This qubit allocation strategy requires a total of $(n-1)\lfloor\log n\rfloor$ qubits for the largest partition $Y_{n}$ since only n1n-1 coefficients are variable and require qubit representation. When the qubit count of the PSA is set by boundary constraints derived from an LLL-reduced basis, it needs $n(n+1)$ qubits, significantly more than IPSA requires. In the case of using constraints from an HKZ-reduced basis, PSA exhibits a comparable $\mathcal{O}(n\log n)$ scaling to IPSA. However, the HKZ-based boundary analysis relies on a classical preprocessing routine with exponential-time complexity, which is often computationally prohibitive to implement. A visual comparison of qubit requirements is shown in Fig. 2
Although theoretical analyses, such as those boundary constraint methods based on LLL and HKZ basis properties to ensure inclusion of the shortest vector, provide valuable upper bounds and specific allocation schemes [25, 26], empirical validation and heuristic insights remain crucial. These are essential for developing qubit allocation strategies that optimize the trade-off between solution quality and qubit count for practical VQA-based SVP solvers like IPSA, especially when aiming to minimize qubit usage beyond worst-case theoretical guarantees.
IV  Numerical Simulations and Results
In this section, we present a systematic numerical comparison of our proposed IPSA against two representative existing methods: the PSA [31] and the IQOAP [28]. All simulations were performed using the Qiskit framework. In Section IV.1, we describe the two distinct sets of SVP instances generated for this study and detail the complete experimental setup for all algorithms. Subsequent sections provide a quantitative analysis across various metrics, confirming the superiority of IPSA.
IV.1  Instances Generation and Experimental Setup
We generated two distinct sets of SVP instances, hereafter referred to as the Benchmark Set and the LLL-Challenging Set, totaling 800 instances across different dimensions. The Benchmark Set comprises 600 instances (200 for each dimension $n\in\{4,5,6\}$) created using a standard procedure [28, 30]: applying random unimodular transformations to the reduced lattice basis. This approach is known to produce instances with difficulty equivalent to that of random lattices [43, 30]. The LLL-Challenging Set consists of an additional 200 instances engineered for $n=6$. These instances were selected by applying the classical LLL algorithm to a large number of randomly generated lattices and retaining only those for which LLL failed to find the true shortest vector, instead finding a different short vector whose length is, on average, only $\approx 2\%$ longer. The ability to solve such instances critically tests an algorithm’s capacity to identify the true shortest vector in the presence of other very short lattice vectors, a scenario where we hypothesize our proposed algorithm offers a distinct advantage.
For coefficient encoding, both IQOAP and our IPSA employ two qubits per variable coefficient. This choice is consistent with the original IQOAP study [28] and satisfies the $\lfloor\log n\rfloor$ qubit requirement for IPSA in this dimensional range. Since the original PSA proposal [31] does not specify a qubit allocation strategy, we benchmarked several variants, denoted as $k$-PSA, where each coefficient is encoded with $k\in\{3,4,5\}$ qubits. However, the computational cost of classically simulating instances requiring more than 22 qubits was prohibitive. Consequently, we had to exclude the 5-PSA variant for $n=5$ and both the 4-PSA and 5-PSA variants for $n=6$.
To evaluate the choice of PQC within our framework, we also implemented an IPSA variant using the QAOA ansatz (IPSA-QAOA) for comparison against the standard HEA-based IPSA. A direct comparison between different PQCs is complex, involving a trade-off between circuit depth, parameter count, and expressivity. To ensure the QAOA ansatz with sufficient expressivity for solving SVP while maintaining a manageable circuit depth, we configured IPSA-QAOA with $p=4$ layers. In contrast, the standard IPSA-HEA was configured with $p=2$ layers. This configuration establishes a meaningful basis for comparing different PQCs for SVP. The key parameters for all algorithms, including the number of layers $p$ in the PQC, are summarized in Table 1.
Table 1:  Parameter configurations for the simulated algorithms. Here, $n$ is the SVP dimension, and $p$ is the number of layers in the PQC. The settings for IQOAP, including a single-layer QAOA ansatz with constrained parameters, are chosen to be consistent with the original study [28].
Algorithms $n$ Max. Qubits $p$ PQC
IPSA 4 6 2 HEA
5 8
6 10
IPSA-QAOA 4 6 4 QAOA
5 8
6 10
IQOAP 4 8 1 QAOA($\beta=\gamma$)
5 10
6 12
3-PSA 4 11 2 HEA
5 14
6 17
4-PSA 4 15 2 HEA
5 19
5-PSA 4 19 2 HEA
For the classical optimization component of all algorithms, we employed the minimize function from the SciPy.optimize library, selecting Powell’s conjugate direction method as the optimizer. The initial parameters for the rotational gates in the PQC were randomly initialized from a uniform distribution over the interval $0,\pi]$. We adopted the default termination criteria from the SciPy implementation: the optimization process stops when the fractional tolerance in either the parameters (xtol) or the cost function value (ftol) is below $1\times 10^{-4}$. All simulations were performed on the Qiskit framework using the StatevectorEstimator or StatevectorSampler.
Figure 3:  Comparative analysis of IPSA-HEA (blue) and IPSA-QAOA (orange) on the Benchmark Set for dimensions $n\in\{4,5,6\}$. The four panels show (a) Success Rate (SR), (b) Approximation Ratio (AR), (c) Total Circuit Depth ($D_{\text{total}}$), and (d) Total CNOT Count ($C_{\text{total}}$). The y-axis values for panels (c) and (d) are presented in units of $10^{6}$. In all plots, the shaded areas represent the interquartile range (25th to 75th percentiles), and the cross markers indicate the median values.
Throughout the following sections, we assess algorithm effectiveness using four key metrics:
  1. 1.
    Success Rate (SR): The fraction of instances where the algorithm successfully finds the shortest vector, defined as $SR=N_{\text{succ}}/N_{\text{total}}$.
  2. 2.
    Approximation Ratio (AR): The quality of the found solution for a given instance, defined as $AR=\lambda_{1}(\mathcal{L})/\|v_{\text{alg}}\|$, where $\lambda_{1}(\mathcal{L})$ is the length of the shortest vector and $\|v_{\text{alg}}\|$ is the length of the vector found by the algorithm. An AR value ranges from 0 to 1, with a value closer to 1 indicating a higher-quality solution. The Average Approximation Ratio (AAR) is then calculated for each algorithm and dimension by taking the mean of the AR values over the 200 corresponding instances.
  3. 3.
    Total Circuit Depth ($D_{\text{total}}$): The cumulative circuit depth across all iterations of the algorithm, $D_{\text{total}}=\sum_{i=1}^{I}d_{i}$, where $d_{i}$ is the depth of the circuit in iteration $i$. This metric reflects, to some extent, the total execution time on quantum hardware.
  4. 4.
    Total CNOT Count ($C_{\text{total}}$): The cumulative number of two-qubit CNOT gates used across all iterations, $C_{\text{total}}=\sum_{i=1}^{I}c_{i}$. This metric quantifies the primary source of noise and error in many NISQ devices.
IV.2  Comparative Analysis: HEA versus QAOA in IPSA
We begin by empirically validating our choice of PQC architecture. This section involves comparing the relative effectiveness of the IPSA implementation using a 2-layer HEA against the version using a 4-layer QAOA. The comparison was conducted on the 600 instances of the Benchmark Set, and the results are presented in Fig. 3.
The figure shows a clear superiority of the HEA implementation across all dimensions. For instance, at $n=6$, the HEA variant achieved an SR of 0.95, whereas the QAOA variant’s SR was only 0.025. The AAR for IPSA-HEA also remained high and close to 1. Its AAR value for $n=6$ was 0.9835. In contrast, the AAR for IPSA-QAOA degraded markedly as the problem size increased, falling from 0.6254 at $n=4$ to 0.4179 at $n=6$.
Figure 4:  Comparative analysis of IPSA (blue) and several existing algorithms, including the $k$-PSA variants (3-PSA in purple, 4-PSA in orange, and 5-PSA in green) and IQOAP (red). The comparison is across dimensions $n\in\{4,5,6\}$ on the Benchmark Set. The panels show (a) Success Rate (SR), (b) Approximation Ratio (AR), (c) Total Circuit Depth ($D_{\text{total}}$), and (d) Total CNOT Count ($C_{\text{total}}$). The y-axis values for (c) and (d) are in units of $10^{5}$. In panels (a) and (b), the brown bars labeled “Initial” represent the metrics of the initial basis. For the data points, the error bars indicate the interquartile range (25th to 75th percentile), and the circular markers denote the median values.
Furthermore, the HEA-based approach was substantially more resource-efficient. At $n=6$, $D_{\text{total}}$ for IPSA-HEA was $0.321\times$106. This result is approximately five times lower than the $1.638\times$106 required by IPSA-QAOA. Its $C_{\text{total}}$ showed an even greater disparity. The value was $0.242\times$106 for IPSA-HEA, nearly ten times lower than the $2.344\times$106 for the QAOA variant. Similar resource advantages for the HEA implementation were observed for $n=4$ and $n=5$.
In summary, these results provide strong empirical evidence that for the SVP instances under consideration within the IPSA framework, the HEA offers a clear and comprehensive advantage over the standard QAOA ansatz, delivering superior solution quality with significantly lower resource consumption. Therefore, for all subsequent experiments presented in this paper, the HEA is adopted as the default PQC for IPSA. While we acknowledge that specific QAOA variants might outperform the standard QAOA ansatz, a detailed investigation of these alternatives is beyond the scope of this work and remains a direction for future research.
IV.3  Comparison with Existing Algorithms
We now compare the effectiveness of IPSA against existing algorithms, namely the various $k$-PSA configurations and IQOAP. The evaluation was performed on the 600 instances of the Benchmark Set using the four metrics defined in Section IV.1. The complete results are summarized in Fig. 4.
Regarding the primary goal of finding the shortest vector, IPSA substantially outperforms the other algorithms. As shown in Fig. 4(a), for dimensions $n=4$, 5, and 6, IPSA achieved an SR of 0.995, 0.985, and 0.95, respectively. By comparison, the SR for IQOAP showed a significant decline from 0.26 to 0.01 across the exact dimensions. The $k$-PSA variants yielded comparatively lower success rates, with SR values generally below 0.1.
Figure 5:  Comparative analysis of IPSA (blue), 3-PSA (purple), and IQOAP (red) on the 200 instances of the LLL-Challenging Set ($n=6$). The classical LLL algorithm (yellow-green) is also included for reference. The panels show (a) Success Rate (SR), (b) Approximation Ratio (AR), (c) Total Circuit Depth ($D_{\text{total}}$), and (d) Total CNOT Count ($C_{\text{total}}$). The y-axis values for (c) and (d) are in units of $10^{5}$. By the selection criteria for this instance set, the SR for the LLL algorithm is zero, but its average AR is high at 0.978.
The advantage of IPSA extends to the quality of approximate solutions. The AAR for IPSA remained close to 1, with a value of 0.9835 for $n=6$. While IQOAP also provides approximate solutions, its AAR values of 0.641, 0.396, and 0.294 were significantly lower than those of IPSA. The PSA variants yielded the lowest AAR values, mainly in the 0.3 to 0.5 range. This result suggests that while iterative algorithms like IPSA and IQOAP find better quality solutions than the non-iterative PSA, the partitioning strategy within IPSA provides a crucial further enhancement.
The resource requirements present a more varied picture. For $n=4$ and 5, the $D_{\text{total}}$ of IPSA was comparable to or lower than that of IQOAP and the higher-qubit $k$-PSA variants. For the largest dimension $n=6$, the $D_{\text{total}}$ for IPSA at $0.321\times 10^{6}$ was higher than for 3-PSA and IQOAP. IQOAP’s lower depth, however, is mainly attributable to its termination after a fixed 50 iterations, a number insufficient to reliably solve the problem at this scale. In terms of the $C_{\text{total}}$, IPSA consistently required fewer gates than IQOAP but more than the PSA variants. This increased resource consumption is a trade-off for the substantial gains in success rate and solution quality.
In summary, the results indicate that both PSA and IQOAP find the shortest vector with low probability, making them unreliable for this task. This low reliability stems from PSA’s tendency to converge on local minima and IQOAP’s propensity for ineffective iterative refinements that fail to improve solution quality. Their limited success at a modest dimension of $n=6$ suggests poor scalability for higher-dimensional problems. In contrast, IPSA achieves the correct solution with high probability. Its integration of iterative refinement and a novel partitioning strategy demonstrates a more effective path to achieving high-quality solutions for the SVP.
IV.4  Effectiveness on Instances with Close Suboptimal Solutions
We then assessed IPSA’s effectiveness on the LLL-Challenging Set. The goal was to evaluate its ability to distinguish the shortest vector from close suboptimal solutions, which is a key difficulty in solving the SVP. These 200 6-dimensional instances from the LLL-Challenging Set are characterized by having LLL-reduced solutions that are, on average, only about 2% longer than the shortest vector length. We compared IPSA against 3-PSA and IQOAP, with the results shown in Fig. 5.
As illustrated in the figure, IPSA maintained a high level of solution quality even in the presence of strong suboptimal attractors. It achieved an SR of 0.86. This result was substantially higher than the SR of 0.165 for 3-PSA and 0.075 for IQOAP. In terms of solution quality, IPSA reached an AAR of 0.996. This result outperforms the AAR values from 3-PSA at 0.869, IQOAP at 0.816, and even the classical LLL algorithm’s initial result of 0.978 in these instances. It is noteworthy that the AAR for 3-PSA and IQOAP improved on this set compared with the $n=6$ Benchmark Set. This phenomenon is because the primary goal of the LLL-Challenging Set was to test an algorithm’s discrimination of close suboptimal solutions. Therefore the instances were generated without the randomizing unimodular transformations used for the Benchmark Set. However, the large error bars associated with both 3-PSA and IQOAP in Fig. 5(b) indicate significant instability in their results across the instance set. This outcome highlights IPSA’s ability to maintain high solution fidelity, even when the solution space contains a prominent local minimum corresponding to the LLL-reduced vector.
This level of solution quality did not require a significant increase in computational overhead compared with the simpler methods. The $D_{\text{total}}$ for IPSA was $1.556\times 10^{5}$, which is comparable to the value for 3-PSA at $1.476\times 10^{5}$. The $C_{\text{total}}$ for IPSA was $1.177\times 10^{5}$, a value also comparable to the $1.269\times 10^{5}$ for 3-PSA. Both of IPSA’s resource metrics were considerably lower than those for IQOAP.
The ability to distinguish between vectors of very similar lengths is a critical hurdle for VQA-based SVP solvers. Heuristic methods are susceptible to being trapped in local minima corresponding to these near-shortest vectors. This issue is particularly pronounced in the LLL-Challenging Set, where the LLL solution and the actual shortest vector reside in similar search regions. IPSA’s 1-tailed partitioning strategy is designed to address this challenge. By design, the strategy separates such closely matched candidate vectors into different search stages or ensures they are differentiated through the deterministic basis update and sorting mechanism of the algorithm. This transformation of a difficult heuristic search into a sequence of more defined sub-problems is central to IPSA’s effectiveness. It allows the algorithm to collect multiple distinct candidate solutions in different iterations, enabling the identification of the global optimum through a final, deterministic comparison of their vector lengths.
In summary, the effectiveness of IPSA on the LLL-Challenging Set highlights the efficacy of its 1-tailed search spaces strategy. This design feature equips the algorithm to perform well in scenarios with substantial suboptimal solution interference, a valuable capability for solving the SVP and other challenging combinatorial optimization problems.
V  Conclusion
In this work, we proposed the IPSA, a “second-generation” iterative VQA-based SVP solver that inherits the ideas of the PSA and the IQOAP framework. It is designed to overcome their fundamental drawbacks, such as ineffective iterations, oversized search spaces, and low circuit efficiency. Its core features, the 1-tailed search space and a stack-based iterative framework, are the key components for achieving this goal. Although PSA and IQOAP may not be the best approach for solving SVP with VQA today, the improved and refined IPSA presented in this work stands as a competitive alternative by mitigating the known drawbacks systematically, offering its own unique merits. Furthermore, IPSA can be employed as a subroutine within algorithms like block Korkin-Zolotarev or be applied to solving problems such as learning with errors, providing an efficient and reliable core component for these complex tasks.
Acknowledgements.
This work is supported by National Natural Science Foundation of China (Grant Nos. 62372048, 62272056, 62371069), and National Key Laboratory of Security Communication Foundation (2025, 6142103042503).